OpenStudy (anonymous):

Find a Cartesian equation relating x and y corresponding to the parametric equations x=5sin(5 t) y=9cos(5 t) Write your answer in the form P(x,y)=0 where P(x,y) is a polynomial in x and y such that the coefficient of y^2 is 25.

OpenStudy (anonymous):

x^2+(25y^2/81)-1 this is the answer I got :/

OpenStudy (amistre64):

it does seem rather elliptic doesnt it ...

OpenStudy (anonymous):

Thats what I did, I used the ellipse formula but I don't know where I'm going wrong.

OpenStudy (amistre64):

r cos(t) = 5 sin(5t) r sin(t) = 9 cos(5t) \[r^2 = \sqrt{25sin^2(5t)+81cos^2(5t)}\] hmmm

OpenStudy (amistre64):

got a spurious ^2 on that last r

OpenStudy (anonymous):

wait where did you get the r, I thought I just had to get "rid" of my t variable by squaring both side and using the unit circle formula,

OpenStudy (amistre64):

just trying to recall a few things is all

OpenStudy (amistre64):

if we unparameterize it, we get x = 5 sin(5t) arcsin(x/5)/5 = t y = 9 cos(5 arcsin(x/5)/5) y = 9 cos(arcsin(x/5))

OpenStudy (amistre64):